Calculate the amount of heat released (kJ) in the complete oxidation of 6.95 g of Al metal at standard condition.
1
Expert's answer
2020-02-04T06:35:07-0500

The standard enthalpy change of formation of aluminum oxide (heat per mole that the substance tranfers to the environment during the process of its formation): "\Delta H^o_f=-1676\text{ kJ/mol}."

1) Write the reaction:


"4\text{Al}+6\text{O}_2\rightarrow2\text{Al}_2\text{O}_3."

We notice that 2 Al atoms are required to produce 1 molecule of aluminum oxide. One molecule of the oxide produce -1676 kJ of heat per mole.

2) How many moles of aluminum reacted?


"n_\text{Al}=\frac{m_\text{Al}}{M_\text{Al}}=\frac{6.95}{27}=0.257\text{ mol}."


3) How many moles of aluminum oxide formed in the reaction? Two times less:


"n_{\text{Al}_2\text{O}_3}=\frac{n_\text{Al}}{2}=\frac{0.257}{2}=0.129\text{ mol}."

4) How much heat the process generated?


"Q=\Delta H\cdot n=-1676\cdot0.129=-216\text{ kJ}."